OpenStudy (anonymous):

Given the linear equation y=7/5x+2 , find the y-coordinates of the points (-5, ), (0, ), and (5, ). Please show all of your work.

OpenStudy (anonymous):

plug in x and find y

OpenStudy (anonymous):

7/5*-5+2=-5 7/5*0+2=2 7/5*5+2=12 (-5,-5)(0,2)(5,12)

OpenStudy (anonymous):

that is what i came up with

OpenStudy (anonymous):

first is correct

OpenStudy (anonymous):

cool, all are correct

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome